OpenStudy (anonymous):

Amber has opened two credit cards in the last year. Use the information on each credit card located in the table below to answer the question that follows. Account Name Account Number Date Opened Balance Past Due Account Status Credit Limit Credit Card A xxxx 5/8/10 $2,397.67 $0.00 PAYS AS AGREED $16,000 Credit Card B xxxx 6/30/10 $1,218.30 $0.00 PAYS AS AGREED $5,500 What is Amber’s overall debt-to-credit ratio?

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

if you can wrap your mind around this one, you have to be god lol

jimthompson5910 (jim_thompson5910):

speaking of wrapping...

OpenStudy (anonymous):

what about wrapping?

jimthompson5910 (jim_thompson5910):

the table/text is

OpenStudy (anonymous):

i know...

OpenStudy (anonymous):

i guess just tell me how to find out a debt to credit ratio?

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

its just this and one more question, then im done!!! for tonight

OpenStudy (anonymous):

@jim_thompson5910 JIMMMMMMMMMMMMMMM, I NEEEEEEEEED YOU FOR TWO MORE PROBLEMS!!!!!!!

jimthompson5910 (jim_thompson5910):

working on it...

OpenStudy (anonymous):

thanks :P

jimthompson5910 (jim_thompson5910):

looking up a term

OpenStudy (anonymous):

what term?

jimthompson5910 (jim_thompson5910):

looking up how exactly debt to credit is defined, I'm getting conflicting definitions at the moment

OpenStudy (anonymous):

bummer :(

jimthompson5910 (jim_thompson5910):

I'll just pick one and go with it

OpenStudy (anonymous):

haha okay, sounds cool with me

jimthompson5910 (jim_thompson5910):

it looks like 2397.67+1218.30 = 3,615.97 is the total balance (for the two cards combined)

jimthompson5910 (jim_thompson5910):

the total credit limit is 16000+5500 = 21,500

OpenStudy (anonymous):

yes, you are right sir

jimthompson5910 (jim_thompson5910):

so the debt to credit ratio is (amount of debt)/(total credit limit) = 3,615.97/21,500 = 0.1681846511628

jimthompson5910 (jim_thompson5910):

so roughly 16.8% of the total credit limit is already used up and tied to other previous debts

OpenStudy (anonymous):

which rounds to .17

jimthompson5910 (jim_thompson5910):

yeah

jimthompson5910 (jim_thompson5910):

so 17%

OpenStudy (anonymous):

which would be 17%

jimthompson5910 (jim_thompson5910):

yep

OpenStudy (anonymous):

yep, thats one of the choices!!!

OpenStudy (anonymous):

can i just ask one more question on here?

jimthompson5910 (jim_thompson5910):

ok gereat

jimthompson5910 (jim_thompson5910):

great*

jimthompson5910 (jim_thompson5910):

yeah sure

OpenStudy (anonymous):

Zoe purchased an entertainment center for $3,257 using a 12-month deferred payment plan with an interest rate of 28.05%. She did not make any payments during the deferment period. What will the total cost of the entertainment center be if she must pay it off within four years after the deferment period?

OpenStudy (anonymous):

the choices are $3,257.00 $4,297.69 $7,195.68 $8,994.60

jimthompson5910 (jim_thompson5910):

the 12 month deferment plan means that no monthly payments are made for 12 months

OpenStudy (anonymous):

i understand

jimthompson5910 (jim_thompson5910):

so it's like a "buy now, pay later" plan

OpenStudy (anonymous):

yep yep

jimthompson5910 (jim_thompson5910):

this has a catch though, just because you're not making payments doesn't mean that interest isn't going up

OpenStudy (anonymous):

i know that too, it sucks

jimthompson5910 (jim_thompson5910):

yeah that's how they get you lol so that initial balance of 3,257 sits in the account for 12 months and it jumps to 3,257(1+0.2805/12)^(12*1) = 4,297.69

OpenStudy (anonymous):

I KNEW IT YAY, i got that exact answer

jimthompson5910 (jim_thompson5910):

so when you finally start that first payment, the balance isn't 3,257..it's really 4,297.69

jimthompson5910 (jim_thompson5910):

we're not done yet

jimthompson5910 (jim_thompson5910):

they put that answer to trick you

OpenStudy (anonymous):

awww. i was so happy

OpenStudy (anonymous):

im sad now... i thought i got it right

jimthompson5910 (jim_thompson5910):

lol finance does that to people...pops a lot of bubbles

OpenStudy (anonymous):

okay, continue then lol

OpenStudy (anonymous):

this is about to be a lot of calculations...

jimthompson5910 (jim_thompson5910):

anyways, we have an initial balance of 4,297.69 and we use this to find the monthly payment needed to pay it all off in 4 years (48 months) M = L[c(1+c)^n]/[(1+c)^n - 1] L = 4297.69 c = 0.2805/12 = 0.023375 n = 48 M = L[c(1+c)^n]/[(1+c)^n - 1] M = 4297.69[0.023375(1+0.023375)^48]/[(1+0.023375)^48 - 1] M = 149.91

jimthompson5910 (jim_thompson5910):

the monthly payment is roughly $149.91 a month for 48 months so you pay a total of 149.91*48 = 7,195.68 dollars

OpenStudy (anonymous):

is that the final answer?!?! or is it just a trick?!?!?!

jimthompson5910 (jim_thompson5910):

why would i post it if it's a trick?

OpenStudy (anonymous):

im just saying haha

jimthompson5910 (jim_thompson5910):

the first one was a trick because a lot of people just stop there and they were expecting you to do so

OpenStudy (anonymous):

thanks @jim_thompson5910 im done for the night!!! Ill def speak with you soon!!! night

jimthompson5910 (jim_thompson5910):

cya